# Dynamical Moiré Systems in Twisted Bilayer Optical Lattices

### v3

@@ Abstract
Moiré related physics in twisted bilayer two-dimensional (2D) materials has attracted widespread interest in condensed matter physics. Simulation of moiré related physics in cold atom platform is expected to outperform the 2D materials thanks to its advantage of higher tunablility. Here, we point out, the cold atom platform enables a new mechanism of moiré lattice formation, with intrinsic “dynamical” character arising from interlayer interaction, in contrast to conventional moiré lattice induced by “static” ways such as single-particle interlayer tunneling or external controls. Specifically, we consider a twisted bilayer Bose-Hubbard model with vanishing interlayer tunneling, and the bilayer is solely coupled through interlayer interaction that originates from contact interaction of atoms. We find that this system hosts a plethora of novel phases unique to this dynamical lattice, including a variety of Mott insulator (MI) and superfluid (SF) phases either preserving or breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, “interlocked” MI, and self-localized phases at commensurate twist angles, which exhibits the characteristics of Bose glass and quasi-many-body localization in the absence of (quasi)disorder or quasicrystalline lattices. Our prediction can be readily observed in current experimental setup of twisted bilayer optical lattices, opening up new avenues for exploring the rich physics of dynamical moiré systems in cold atoms.
Introduction.—Moiré physics has been extensively studied in the field of condensed matter physics in recent years. One prominent feature that stimulates tremendous interest is the emergence of flat band by simply tuning the twist angle of bilayer two-dimensional (2D) materials, which offers an intriguing platform for exploring phases that electron-electron interaction plays a significant role, such as correlated insulators Cao et al. (2018a); Li et al. (2021a), superconductivity Cao et al. (2018b), magnetism Gong et al. (2017); Huang et al. (2017); Chen et al. (2020) and generalized Wigner crystals Regan et al. (2020); Li et al. (2021b). On the other hand, topological phases also emerge in moiré superlattices, e.g., quantum anomalous Hall insulators (QAHI) Serlin et al. (2020); Li et al. (2021c); Zhou et al. (2021), $\mathbb{Z}_{2}$ topological insulators Wu et al. (2019); Zhao et al. (2024), high-order topological insulators Park et al. (2019), and fractional QAHI Cai et al. (2023); Park et al. (2023); Zeng et al. (2023); Xu et al. (2023); Lu et al. (2024). The rich many-body and topological physics observed in bilayer 2D materials have clearly demonstrated that twisting serves as a novel and effective means for manipulating material properties and unlocking diverse and fascinating quantum phenomena Kennes et al. (2021).
@@
So far, the moiré lattices observed in both 2D materials and cold atoms share a common feature: they are “static”, in the sense that their forms are uniquely fixed by deterministic factors such as interlayer tunneling, structural relaxations, or external controls like dielectric screening Bistritzer and MacDonald (2011); Wang et al. (2017); Ruiz-Tijerina and Fal’ko (2019); Li et al. (2021d); Xu et al. (2021); Kim et al. (2024); Zhang et al. (2024); Gu et al. (2024); He et al. (2024). In this work, we propose a distinct mechanism of realizing moiré lattice solely induced by particle-particle interaction from two layers or pseudospins, which we refer to as “dynamical” moiré lattice  [^1] This is to be contrasted with dynamical control of moiré lattice already realized in 2D materials, where the moiré potential in the target layer is uniquely imprinted by substrate layers and has no feedback on the substrate layers Kim et al. (2024); Zhang et al. (2024); Gu et al. (2024); He et al. (2024). For each layer/pseudospin of lattice, the effect of the other layer is brought by particle-particle interaction, which relies on particle density distribution, an intrinsic dynamical variable. Therefore, each layer can feel dynamical feedback from the other layer and the ultimate density distribution is a compromise between the two layers. Realizing such a dynamical moiré system in twisted 2D materials is challenging, but it can be easily achieved in current cold atom platforms by simply turning off interspin coupling in twisted spin-dependent lattices Meng et al. (2023).
@@
Figure 1: (a) Schematic illustration of the dynamical moiré system. (b) Sketch of a twisted bilayer optical lattice at the twisted angle $\theta=2\arctan(1/3)=36.87^{\circ}$. Red (blue) triangles correspond to lattice sites of $\uparrow$ ($\downarrow$) layer. The purple circle denotes the external cylinder box trap. (c) Schematic illustration of the origin of the interlayer interaction. Filled triangles denote the lattice sites occupied by atoms. The overlap between atomic wave functions (denoted by the translucent circles) of atoms on different layers gives rise to the interlayer interaction. (d) Zoo of rich phases in the dynamical moiré system. The left (right) column of each subplot corresponds to the ($\uparrow$($\downarrow$)↑ ( ↓ )-layer, respectively. From left to right, up to down, ($\theta$, $\rho_{\uparrow}$, $\rho_{\downarrow}$, $g_{\uparrow\downarrow}/g_{\sigma\sigma}$, $J/U(g_{\sigma\sigma})$) assume the values ($7.63^{\circ}$, $3/5$, $3/5$, $0.5$, $0.06$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.9$, $0.048$), ($36.87^{\circ}$, $7/10$, $4/5$, $0.9$, $0.04$), ($22.62^{\circ}$, $2/5$, $3/5$, $0.9$, $0.024$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.3$, $0.008$), and ($7.63^{\circ}$, $7/10$, $3/5$, $0.5$, $0.012$), respectively. See text for more details.
@@
Specifically, we consider two species of ultracold atoms loaded in a twisted bilayer spin-dependent optical lattice and find that, an emergent moiré lattice from particle-interaction can form, hosting much more rich phases than static lattice counterpart (see Fig. 1), including phases with and without moiré lattice symmetry and Mott insulator (MI) phases induced by interlayer interaction even for partial filling. More interestingly, for total unit filling, the ground state can become an “interlocked” MI with spontaneously formed moiré lattice. For a general total filling, the Bose glass phase exists, suggesting that the ground state tends to localize due to the interplay between twisting and interlayer interaction, even at commensurate twist angles.
@@
System and model.—Motivated by the recent realization of the twisted bilayer optical lattices Meng et al. (2023), we consider a closely related but simpler experimental set-up, namely, a twisted bilayer optical lattice system without the additional microwave field for realizing the interlayer coupling in the recent experiments Meng et al. (2023). More specifically, the system consists of two species of bosonic atoms that correspond to two hyperfine states (for instance, ${}^{87}\mathrm{Rb}$ atoms in hyperfine states $|F=1,m_{F}=1\rangle$ and $|F=2,m_{F}=0\rangle$ considered in Ref. Meng et al. (2023)), with each species loaded in one layer of a spin-dependent twisted bilayer square optical lattice, respectively [see Fig. 1(a)].
@@
The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation (see Supplemental Material (SM) Sup for a detailed derivation),
@@
The third part of $\hat{H}$ captures the distinct way that the twist imposes its physical influences even without the interlayer tunneling. It originates from the finite overlaps between wave functions of atoms on different layers with their inter-species contact interaction being $g_{\uparrow\downarrow}$ [see Fig. 1(c)], and consists of interlayer interaction terms $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$, with the interaction strength $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ assuming the twist angle $\theta$ dependence. This dependence reflects how the overlaps between Wannier wave functions of different sites on different layers are changed when one tunes $\theta$.
Figure 2: Phase diagrams and typical configurations at the balanced and imbalanced fillings. (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the balanced filling with $\rho_{\uparrow}=4/5$ and $\rho_{\downarrow}=4/5$ (twist angle is fixed at $\theta=36.87^{\circ}$). (b) Similar to (a), but at the imbalanced filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=4/5$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively. See text for more details.
@@
We investigate this system by calculating its ground state via the bosonic Gutzwiller variational approach Krauth et al. (1992); Jaksch et al. (1998); Lanatà et al. (2012) (see SM Sup for technical details). We have chosen the twist angle $\theta=2\arctan(\bar{m}/\bar{n})$ with $\bar{m},\bar{n}$ being two natural numbers Wang et al. (2020); Meng et al. (2023), i.e., the commensurate condition for twisted square lattice, and focus on the case with $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}$. If not specified otherwise, we perform the numerical simulation on a $37\times 37$ twisted bilayer optical lattice with an external 2D cylinder box trap Navon et al. (2021) whose radius is 18 lattice constants (denoted by the solid circles in the figures). In the following, we focus on the most interesting case of fractional filling for both layers, and reserve the results of integer filling in the Supplemental Material Sup (the impact of interlayer interaction is not as dramatic in the latter case).
@@
Before presenting the numerical results, we first notice that in conventional moiré systems Bistritzer and MacDonald (2011); Wang et al. (2017); Ruiz-Tijerina and Fal’ko (2019); Meng et al. (2023), one layer of lattice actually influences the physics on the other layer in a “static” way, in the sense that the influences are solely determined by the configuration of the twisted bilayer lattice through the single-particle interlayer tunneling. In sharp contrast, here, these influences are “dynamical”, in the sense that one layer of lattice impose its influences on the physics of the other via the dynamical variables on this layer (particle number operators in this case), as clearly manifested by the interlayer interaction terms of the Hamiltonian (1), giving rise to a new type of moiré system we dubbed as “dynamical moiré system”. This intrinsic dynamical character could make the system feature not only similar phases in the conventional moiré system, but also many intriguing phases without counterpart in the latter [see Fig. 1(d) for an overview].
Conventional moiré physics without interlayer tunneling.—In conventional bosonic moiré systems, the interlayer tunneling causes layer hybridization of single particle orbitals, and thus particle filling on the hybridized orbitals results in phases with two layer sharing the same character. Indeed, as observed in recent experiments Meng et al. (2023), it can accommodate SF-SF and MI-MI (denoting SF/MI phase in $\uparrow$/$\downarrow$ / ↓ layer), with the particle density assuming the moiré lattice symmetry. In fact, these observations can be easily explained by treating the system as if the atoms on one layer experience an additional static potential from the other layer of the lattice Meng et al. (2023). Speculating along this line, we notice that although there is no interlayer tunneling in the dynamical moiré system studied here, atoms residing on one layer can nevertheless feel an additional potential provided by the atoms residing on the other layer via the interlayer interactions $\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$. This indicates the system may accommodate similar conventional moiré physics even without the interlayer tunneling.
In Fig. 2(a), we map out the phase diagram of the system in a balanced filling case with $\rho_{\uparrow}=\rho_{\downarrow}=4/5$ ($\rho_{\sigma}\equiv\langle\sum_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{i}_{\sigma}}\rangle/N_{\sigma}^{\mathrm{lat}}$ with $N_{\sigma}^{\mathrm{lat}}$ being the number of lattice sites of the $\sigma$-layer within the cylinder trap). In the relatively large hopping regime, we can see that the system indeed support SF-SF phase (denoted as SF in the phase diagrams) similar to conventional static moiré systems, with both layers manifesting the SF order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ assuming the moiré lattice symmetry [see the first plot in Fig. 2(c)]. While in the relatively small hopping regime, we notice that despite the filling factor on each layer being non-integer, increasing the inter-species interaction strength $g_{\uparrow\downarrow}$ can drive the system into the induced MI phase characterized by regions with integer-filled particles and vanishing SF order parameter in either layer [see the last plot in Fig. 2(c)], including MI-SF, MI-MI and SF-MI phases. This is in sharp contrast to the single-component Bose gases in optical lattices, where the MI can only exist at integer fillings.
The induced MI phase occurs because increased $g_{\uparrow\downarrow}$ enhances the effective potential experienced by atoms on one layer imposed by atoms on the other layer, causing localization. This scenario is reminiscent of conventional bosonic moiré systems, where increasing the interlayer tunneling makes one layer impose a stronger additional static lattice potential on the other layer, leading to stronger localization Meng et al. (2023). However, it’s important to recognize a key difference here: the effective potential is a dynamical one determined by the dynamical variables, and hence does not necessarily share the same spatial symmetry of the underlying optical lattice. For balanced filling case, indeed, we have not found MI phase with both layer assumes the moiré lattice symmetry. Instead, as shown in the second and third plot in Fig. 2(c), we find this dynamical moiré system can manifest rich discrete rotational symmetric MI and SF phases with no counterparts in the conventional bosonic moiré systems. However, away from the balanced filling, we indeed find the system can support MI phase with density distributions on both layers assuming the moiré lattice symmetry. For instance, as shown in the last two plots in Fig. 2(d), this kind of moiré MI phase can emerge in the induced MI regime of the system at an imbalanced filling $\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5$.
Although the dynamical moiré system can manifest SF or MI phase similar to the ones found in conventional bosonic moiré systems Meng et al. (2023), there are fundamental differences. Instead of being coupled by interlayer tunneling, the atoms on different layers are coupled by the interlayer density-density interactions. Consequentially, the total particle number of each layer is conserved and can be different. These fundamental differences thus naturally indicate much richer physics with no counterparts in conventional bosonic moiré systems can emerge, as we shall discuss below.
Beyond the conventional moiré physics.—By comparing the physics at balanced and imbalanced filling as showcased in Fig. 2, one clearly notices the strong physical influences of the filling difference. Since the total particle number of each layer is conserved separately, the filling difference between the two layers in fact provides an extra tuning knob of new physics compared with the conventional bosonic moiré systems. In Fig. 3(a), we systematically investigate the influences of the filling difference by keeping the filling factor of the $\downarrow$-layer fixed at $\rho_{\downarrow}=4/5$ and changing the filling factor of the $\uparrow$-layer $\rho_{\uparrow}$ from $1/10$ all the way to $4/5$. As we can see from the first plot in Fig. 3(c), when the filling of the $\uparrow$-layer is small ($\rho_{\uparrow}=1/10$), the $\downarrow$-layer is in a SF phase and it imposes strong influences on atoms on the $\uparrow$-layer by localizing them and making them in induced MI phase despite $\rho_{\uparrow}=1/10$ being non-integer. As the $\rho_{\uparrow}$ gradually increases [see the last two plots in Fig. 3(c)], the $\uparrow$-layer also imposes an enhanced localization effects on the atoms on the $\downarrow$-layer by driving a SF-MI transition on this layer. We also observe that the average localization effect experienced by each atom in the $\uparrow$-layer decreases, making the $\uparrow$-layer transition into a SF phase at relatively high filling with $\rho_{\uparrow}=1/2$. This phenomenon is reminiscent of what happens in high-temperature superconductors, where doping can lead to significant changes in the system’s transport properties Keimer et al. (2015). More interestingly, for the total unit filling, the system hosts a kind of “interlocked” MI with “interlocked” density distributions [see the second plot of Fig. 3(c), the lower left plot of Fig. 1(d) and the fourth plot of Fig. 2(d)].
Similar to the conventional moiré system, the twist angle also has a significant impact here, as clearly shown in Fig. 3(b). For the case at the filling $(\rho_{\uparrow}=1/10,\rho_{\downarrow}=3/5)$, in the intermediate hopping regime, we see that at a relatively small twist angle $\theta=7.63^{\circ}$, the system is in the SF-SF phase. As the twist angle increases to $\theta=22.62^{\circ}$, the atoms in the $\uparrow$-layer are localized due to the interaction from the $\downarrow$-layer and enter the induced MI phase, forming the phase MI-SF [see the second plot of Fig. 3(d)]. Similar situation is also found at even larger twist angle, for instance, $\theta=36.87^{\circ}$ [see the last plot of Fig. 3(d)].
Interestingly, we notice that density distributions in the last two plots of Fig. 3(d) assume rather random pattern. In these cases, the atoms on $\downarrow$-layer in fact provide a “disordered” effective potential to the atoms lying on the $\uparrow$-layer. Thus, one naturally expects that this “disordered” effective potential quenches hopping of the $\uparrow$-atoms and makes them localized, which can be interpreted as the interlayer interaction induced “self-localization” between the atoms on the two layers. Among the states with irregular density pattern, such as the third plot of Fig. 3(d) and the last plot of Fig. 1(d), there exist states with disconnected superfluid islands, exhibiting the characteristics of Bose glass. Thus, this system supports Bose glass without extrinsic disorder, in contrast to previous studies with either (quasi)disorder or quasicrystalline lattices Fisher et al. (1989); Fallani et al. (2007); Pasienski et al. (2010); Söyler et al. (2011); Gadway et al. (2011); D’Errico et al. (2014); Meldgin et al. (2016); Johnstone et al. (2021); Ciardi et al. (2023). The presence of Bose glass is associated with the spontaneous breaking of $C_{4}$ rotational symmetry of the system, indicating that there are four degenerate ground states. As the presence of Bose glass at ground state usually implies many-body localization (MBL) at excited states Abanin et al. (2019), this system also probably hosts (quasi-)MBL without extrinsic disorder Schiulaz et al. (2015); Yao et al. (2016); Smith et al. (2017); Karpov et al. (2021). Note that here either the Bose glass or the possible quasi-MBL are found at commensurate twist angles, i.e., the system has translational invariance in the bulk, in contrast to previously investigated localization at incommensurate angles Huang and Liu (2019); Paul et al. (2023); Madroñero et al. (2024) or the interacting 2D Aubry-André model with quasi-periodicity Szabó and Schneider (2020); Štrkalj et al. (2022). The possible quasi-MBL here may be explained by a mechanism similar to the exponentially slow relaxation predicted in certain translation-invariant systems Schiulaz et al. (2015); Yao et al. (2016). Further exploration of the relation between the localization found here with MBL and the generalization to incommensurate twist angles will be an intriguing subject for future study.

### v4

# Interaction induced moiré systems in twisted bilayer optical lattices

@@ Abstract
Moiré related physics in twisted bilayer two-dimensional (2D) materials has attracted widespread interest in condensed matter physics. Simulation of moiré related physics in cold atom platform is expected to outperform the 2D materials thanks to its advantage of higher tunablility. Here, we demonstrate that, the cold atom platform enables a new mechanism of moiré lattice formation, induced by interlayer interaction with intrinsic “dynamical” character, in contrast to conventional moiré lattice induced by “static” ways such as single-particle interlayer tunneling. Specifically, we consider a twisted bilayer Bose-Hubbard model with vanishing interlayer tunneling, and the bilayer is solely coupled through interlayer interaction that originates from contact interaction of atoms. We find that this system hosts a plethora of novel phases unique to this dynamical lattice, including a variety of Mott insulator (MI) and superfluid (SF) phases either preserving or breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, “interlocked” MI, and self-localized phases at commensurate twist angles, which exhibits the characteristics of Bose glass and quasi-many-body localization in the absence of (quasi)disorder or quasicrystalline lattices. Our prediction can be readily observed in current experimental setup of twisted bilayer optical lattices, opening up new avenues for exploring the rich physics of interaction induced moiré systems in cold atoms.
I Introduction
Moiré physics has been extensively studied in the field of condensed matter physics in recent years. One prominent feature that stimulates tremendous interest is the emergence of flat band by simply tuning the twist angle of bilayer two-dimensional (2D) materials, which offers an intriguing platform for exploring phases that electron-electron interaction plays a significant role, such as correlated insulators Cao et al. (2018a); Li et al. (2021a), superconductivity Cao et al. (2018b), magnetism Gong et al. (2017); Huang et al. (2017); Chen et al. (2020) and generalized Wigner crystals Regan et al. (2020); Li et al. (2021b). On the other hand, topological phases also emerge in moiré superlattices, e.g., quantum anomalous Hall insulators (QAHI) Serlin et al. (2020); Li et al. (2021c); Zhou et al. (2021); Qiu et al. (2023), $\mathbb{Z}_{2}$ topological insulators Wu et al. (2019); Zhao et al. (2024), high-order topological insulators Park et al. (2019), and fractional QAHI Cai et al. (2023); Park et al. (2023); Zeng et al. (2023); Xu et al. (2023); Lu et al. (2024). The rich many-body and topological physics observed in bilayer 2D materials have clearly demonstrated that twisting serves as a novel and effective means for manipulating material properties and unlocking diverse and fascinating quantum phenomena Kennes et al. (2021).
@@ I Introduction
So far, the moiré lattices observed in both 2D materials and cold atom studies share a common characteristic: they are “static”, in the sense that their form is uniquely determined by interlayer tunneling, structural relaxations, or external controls such as dielectric screening Bistritzer and MacDonald (2011); Wang et al. (2017); Ruiz-Tijerina and Fal’ko (2019); Li et al. (2021d); Xu et al. (2021); Kim et al. (2024); Zhang et al. (2024); Gu et al. (2024); He et al. (2024). Here in this work, we propose a distinct mechanism of realizing moiré lattice purely induced by particle-particle interaction from two layers or two pseudospins, which we refer to as “dynamical” moiré lattice. With vanishing interlayer coupling at single-particle level, the two layers are completely isolated, and thus there is no moiré lattice at this level. For each layer/pseudospin of lattice, the effect of the other layer is brought by particle-particle interaction, which relies on particle density, a dynamical variable  [^1] Therefore, each layer can feel dynamical feedback from the other layer and the ultimate density distribution is a compromise between the two layers. This is to be contrasted with dynamical control of moiré lattice already realized in 2D materials, where the moiré potential in target layer is uniquely imprinted by substrate layers and has no feedback on the substrate layers Kim et al. (2024); Zhang et al. (2024); Gu et al. (2024); He et al. (2024). Realizing such a dynamical moiré system in twisted 2D materials is challenging, but it can be easily achieved in modern cold atom platforms by simply turning off interspin coupling in the twisted spin-dependent lattices Meng et al. (2023).
@@
Figure 1: (a) Schematic illustration of the dynamical moiré system. (b) Sketch of a twisted bilayer optical lattice at the twisted angle $\theta=2\arctan(1/3)=36.87^{\circ}$. Red (blue) triangles correspond to lattice sites of $\uparrow$ ($\downarrow$) layer. The purple circle denotes the external cylinder box trap. (c) Schematic illustration of the origin of the interlayer interaction. Filled triangles denote the lattice sites occupied by atoms. The overlap between atomic wave functions (denoted by the translucent circles) of atoms on different layers gives rise to the interlayer interaction. (d) Zoo of rich phases in the dynamical moiré system. The left (right) column of each subplot corresponds to the ($\uparrow$($\downarrow$)↑ ( ↓ )-layer, respectively. From left to right, up to down, ($\theta$, $\rho_{\uparrow}$, $\rho_{\downarrow}$, $g_{\uparrow\downarrow}/g_{\sigma\sigma}$, $J/U(g_{\sigma\sigma})$) assume the values ($7.63^{\circ}$, $3/5$, $3/5$, $0.5$, $0.06$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.9$, $0.048$), ($36.87^{\circ}$, $7/10$, $4/5$, $0.9$, $0.04$), ($22.62^{\circ}$, $2/5$, $3/5$, $0.9$, $0.024$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.3$, $0.008$), and ($7.63^{\circ}$, $2/5$, $3/5$, $0.5$, $0.024$), respectively.
@@
Specifically, we consider two species of ultracold atoms loaded in a twisted bilayer optical lattice and find that, an emergent moiré lattice from particle interaction can form, hosting much more rich phases than static lattice counterpart (see Fig. 1), including phases with and without moiré lattice symmetry and Mott insulator (MI) phases induced by interlayer interaction even for partial filling. More interestingly, for total unit filling, the ground state can become an “interlocked” MI with spontaneously formed moiré lattice. For a general total filling, the Bose glass (BG) phase exists, suggesting that the ground state tends to localize due to the interplay between twisting and interlayer interaction, even at commensurate twist angles.
II System and model
@@
Motivated by the recent realization of the twisted bilayer optical lattices Meng et al. (2023), we consider a closely related but simpler experimental setup, namely, a twisted bilayer optical lattice system without the additional microwave field for realizing the interlayer coupling in the recent experiments Meng et al. (2023). More specifically, the system consists of two species of bosonic atoms that correspond to two hyperfine states (for instance, ${}^{87}\mathrm{Rb}$ atoms in hyperfine states $|F=1,m_{F}=1\rangle$ and $|F=2,m_{F}=0\rangle$ considered in Ref. Meng et al. (2023)), with each species loaded in one layer of a spin-dependent twisted bilayer square optical lattice, respectively [see Fig. 1(a)].
@@
The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation [see Appendix A for a detailed derivation],
@@
The third part of $\hat{H}$ captures the distinct way that the twist imposes its physical influences even without the interlayer tunneling. It originates from the finite overlaps between wave functions of atoms on different layers with their inter-species contact interaction being $g_{\uparrow\downarrow}$ [see Fig. 1(c)], and consists of interlayer interaction terms $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$, with the interaction strength $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ assuming the twist angle $\theta$ dependence. As the overlaps of the wave functions depend directly on the spatial distance of sites $\mathbf{i}_{\uparrow}$ and $\mathbf{j}_{\downarrow}$, the interaction strength $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ has explicit dependence on both spatial distance and twist angle $\theta$. While many studies have investigated multicomponent Bose-Hubbard models in the context of ultracold atoms, our study focuses on the largely unexplored many-body physics of two-component ultracold bosons in twisted bilayer optical lattices. The rich physics arising from the twist-dependent interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ in such systems has not yet been systematically studied.
@@
We investigate this system by calculating its ground state via the bosonic Gutzwiller variational approach Krauth et al. (1992); Jaksch et al. (1998); Lanatà et al. (2012) (see Appendix A for technical details). We have chosen the twist angle $\theta=2\arctan(\bar{m}/\bar{n})$ with $\bar{m},\bar{n}$ being two natural numbers Wang et al. (2020); Meng et al. (2023), i.e., the commensurate condition for twisted square lattice, and focus on the case with $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}$. We perform the numerical simulation on a $37\times 37$ twisted bilayer optical lattice with an external 2D cylinder box trap Navon et al. (2021) whose radius is 18 lattice constants (denoted by the solid circles in the figures). We additionally investigate the finite-size effects and find no significant differences between the results here and those on a larger system (see the Appendix C). In the following, we focus on the most interesting case of fractional filling for both layers, and reserve the results of integer filling in the Appendix B (the impact of interlayer interaction is not as dramatic in the latter case).
@@
Before presenting the numerical results, we first notice that in conventional moiré systems Bistritzer and MacDonald (2011); Wang et al. (2017); Ruiz-Tijerina and Fal’ko (2019); Meng et al. (2023); Qiu et al. (2023), the moiré lattices in twisted bilayers are induced by single-particle interlayer tunneling. Particle interactions can be further considered within the framework of these preformed moiré lattices. In sharp contrast, here, the twisted bilayers are completely isolated at single-particle level, and the moiré lattice arises purely from the interlayer interaction terms of the Hamiltonian (1). This novel mechanism of moiré formation could make the system feature not only similar phases in the conventional moiré system, but also many intriguing phases without counterpart in the latter [see Fig. 1(d) for an overview].